1. 7=(?)% of 28 2. 3/16=(?)% x 3/4 3.9is what percent of 27
4. 12:30 as a percentage

Answer by checkley77(12844)

1. 7=(?)% of 28
7/28= 1/4=.25=25% Answer.
2. 3/16=(?)% x 3/4
(3/16)/(3/4)=3/16*4/3=12/8=1/4=.25=25% Answer.
3. 9is what percent of 27
9/27=1/3=.333=33.3% Answer.
4. 12:30 as a percentage
12/30=.4=40% Answer.